OpenStudy (anonymous):

Thanks everyone. I am afraid I don't remember this. If 3x+5 =4, what is the value of 3x-2?

hartnn (hartnn):

find 3x=.... first by subtracting 5 from both sides, what u get ?

hartnn (hartnn):

or you could simply subtract 7 from both sides

OpenStudy (anonymous):

make this 3x+5 =4 in form of 3x - 2 = ?

OpenStudy (anonymous):

so wat shuld u do....

OpenStudy (anonymous):

Is the answer 2

hartnn (hartnn):

nopes

OpenStudy (anonymous):

Sorry, theyears have been cruel.

hartnn (hartnn):

3x+5 = 4 3x = 4-5 = -1 (subtracting 5 from both sides ) 3x-2 = -1-2 (subtracting 2 from both sides ) =-3 got this ?

OpenStudy (anonymous):

Just subtract 7 from both sides and you would directly get what u need. Ya -3 is correct